1) (x-4)^2/9+y^2+4y+4/25=1
Solve for y, find center, focus, name the curve
Would be very thankful for help in this have no clue where to start!
Answer by lwsshak3(11628) 
 (x-4)^2/9+y^2+4y+4/25=1
Solve for y, find center, focus, name the curve
(x-4)^2/9+(y^2+4y+4)/25=1
(x-4)^2/9+(y+2)^2/25=1
This is an equation of an ellipse with vertical major axis of the standard form:
x-h)^2/b^2+(y-k)^2/a^2, a>b, (h,k)=(x,y) coordinates of center
For given equation:(x-4)^2/9+(y+2)^2/25=1
center:(4,-2)
a^2=25
b^2=9
c^2=a^2-b^2=25-9=16
c=√16=4
focus: (4,-2±c)=(4-2±4)=(4,-6) and (4,2)
| |
|